Circular No. 6843

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1998W IN NGC 3075 W. Li, M. Modjaz, R. R. Treffers, and A. V. Filippenko, University of

California at Berkeley, report their discovery of an apparent supernova during the course of the Lick Observatory Supernova Search (cf. IAUC 6627) with the 0.8-m Katzman Automatic Imaging Telescope. SN 1998W was found and confirmed on unfiltered CCD images obtained on Mar. 16.3 (mag ∼ 17.3) and 17.2 UT (mag ∼ 17.4), and it is located at α = 9^{h}58^{m}54^{s}.83, δ = +14^{o}25′25′′.7 (equinox 2000.0), which is ∼ 20′′ west and 12′′ north of the nucleus of NGC 3075. An image of the same field taken on Jan. 25 (limiting mag ∼ 18.5) shows no star at this position.

SAX J1810.8–2609 P. Ubertini and A. Bazzano, Istituto di Astrofisica Spaziale (IAS),

CNR, Rome; J. Heise and J. in ’t Zand, Space Research Organization of the Netherlands, Utrecht; A. Antonelli, BeppoSAX Science Operation Center, Rome; and L. Piro, IAS, report on a BeppoSAX follow-up observation of SAX J1810.8–2609 (cf. IAUC 6838) with the Narrow Field Instruments on Mar. 12: “An improved position has been obtained, corresponding to α = 18^{h}10^{m}44^{s}.5, δ = −26^{o}06′39′′ (equinox 2000.0; error radius 1′); this is 2′.5 from the previously reported position. The average flux for the source in the band 2–10 kev corresponds to 1.5 × 10^{−10} erg cm^{−2} s^{−1} and photon index 2.2.”

SUPERNOVA 1998S IN NGC 3877 Magnitude estimates, visual unless otherwise noted (comparison-star

magnitudes from an AAVSO sequence): Mar. 15.814 UT, 12.29 (M. Bignotti, Lumezzane, Italy; CCD); 15.9, 12.1 (L. Kiss, Szeged, Hungary); 16.85, 12.2 (Kiss); 16.93, 11.8 (Carvajal); 17.146, 12.0 (C. E. Spratt, Victoria, BC); 17.22, 12.5 (D. Borgman, W. Dillon, and K. Rivich, Damon, TX; unfiltered CCD); 17.8, 12.1 (Kiss).

SUPERNOVA 1998T IN IC 694 Visual magnitude estimates: Mar. 15.795 UT, 14.3 (K. Hornoch, Lele-

kovice, Czech Republic; GSC comparison stars); 15.803, 13.9 (M. Plsek, Lelekovice, Czech Republic).

1998 March 18  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green